OpenStudy (anonymous):

Can someone help me solve these word problems? The sum of two numbers is 26 and their difference is 20. Find each of the numbers.

OpenStudy (texaschic101):

x + y = 26 ---> x = 26 - y now sub 26 - y in for x in the other equation x - y = 20 26 - y - y = 20 26 - 2y = 20 -2y = 20 - 26 -2y = - 6 y = -6/-2 y = 3 now sub 3 in for y in either equation to find x x + y = 26 x + 3 = 26 x = 26 - 3 x = 23 check.. x - y = 20 23 - 3 = 20 20 = 20 (correct) check.. x + y = 26 23 + 3 = 26 26 = 26(correct) your numbers are : x = 23 and y = 3